OpenStudy (anonymous):

...Determine whether f(x) and g(x) are inverses of each other. f(x)=(2)/(3)x+6 g(x)=(3)/(2)x-9

OpenStudy (ash2326):

We have \[f(x)=\frac{2}{3} x+6\] \[g(x)=\frac{3}{2} x-9\] If f(x) and g(x) are inverse of each other then \[f(g(x))=x\] Let's check for this \[f(g(x))=\frac{2}{3} g(x)+6\] \[g(x)=\frac{3}{2} x-9\] Now \[f(g(x))=\frac{2}{3} (\frac{3}{2} x-9)+6=x-\frac{2}{3} \times 9+6=x-6+6=x\] Hence f(x) and g(x) are inverse of each other